OpenStudy (anonymous):

24=-8(2n+8)-5n solve each equation.

OpenStudy (anonymous):

24 = -16n - 64 -5n 24 = -21n -64 24 + 64 = -21n 88 = -21n 88/-21 = n n = -88/21

OpenStudy (anonymous):

24=-8(2n+8)-5n 24=-16n-64-5n 24=-21n-64 24+64=-21n 88/-21=n